Suppose there are currently 80 machines, depreciation is 10% and the optimal capital stock for

next year (where MPK=UC) is 92 machines. What should net and gross investment equal in this

case?
Answer #1

Current capital is 80 machines.

Depreciation is 10%. Depreciation = 0.1*80 = 8 machines.

So, the capital remaining after depreciation for the next year = 80 - 0.1*80 = 72 machines.

For the next year, the optimal stock is 92 machines.

So, the gross investment will be 92-72 = 20 machines

The net investment = Gross investment - Depreciation

Net investment = 20 - 8 = 12 machines.

IF YOU FIND THIS ANSWER HELPFUL, PLEASE UP VOTE.